In a right triangle, one of the legs is 2 feet longer than the shorter leg. The hypotenuse is 10 feet. Find the lenghths of the missing 2 sides of the triangle.
Answer by checkley77(12844)

X^2+(X+2)^2=10^2
X^2+X^2+4X+4=100
2X^2+4X+4-100=0
2X^2+4X-96=0
2(X^2+2X-48)=0
2(X+8)(X-6)=0
X-6=0
X=6 FEET FOR THE SHORTER SIDE.
6+2=8 FEET FOR THE LONGER SIDE.
PROOF:
6^2+8^2=10^2
36+64=100
100=100